OpenStudy (anonymous):

Suppose the magnitude of C=6 and the magnitude of D=4, and l C x D l= 12. Find C dot D?

OpenStudy (anonymous):

are these vectors or just regular numbers?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the answer is 12 root 3, but I'm really not sure how to start the problem

OpenStudy (anonymous):

it's been some time since I've done vectors lol give me a minute

OpenStudy (anonymous):

what is l C x D l exactly?

OpenStudy (anonymous):

cross product?

OpenStudy (anonymous):

magnitude of the cross product of C and D?

OpenStudy (anonymous):

alright, busting out my old notes for this lol you know that |C x D| = |C||D|sin(theta) right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you know all the values there except for theta so you can solve for that

OpenStudy (anonymous):

is C dot D= lCllDl cos(theta) also?

OpenStudy (anonymous):

yes, so all you need is the angle

OpenStudy (anonymous):

how do I find the angle?

OpenStudy (anonymous):

|C x D| = |C||D|sin(theta) (12) = (6)(4)sin(theta) (12) = (24)sin(theta) 12/24 = sin(theta) 1/2 = sin(theta)

OpenStudy (anonymous):

so I plugged in the numbers and got 60 degrees

OpenStudy (anonymous):

oh! I got it!

OpenStudy (anonymous):

6*4sin(60)= 12 rt 3. phew

OpenStudy (anonymous):

thank you!!

OpenStudy (anonymous):

yeah no problem, good job!